OpenStudy (anonymous):

The point (1, −1) is on the terminal side of angle θ, in standard position. What are the values of sine, cosine, and tangent of θ? Make sure to show all work.

OpenStudy (anonymous):

@ParthKohli @Preetha @hartnn @jigglypuff314 @KendrickLamar2014 @Lexi_Loves

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

I will fan and write a testimonial.

OpenStudy (lexi_loves):

Use Pythagorean Theorem to find the hypotenuse.

OpenStudy (anonymous):

Ok, hold on.

OpenStudy (anonymous):

But it's just one point, I don't understand.

OpenStudy (lexi_loves):

So r r=12+−12−−−−−−−√

OpenStudy (lexi_loves):

hold on

OpenStudy (anonymous):

r = √1^2 + (-1)^2?

OpenStudy (anonymous):

Is r √2?

OpenStudy (lexi_loves):

yes that is correct :)

OpenStudy (anonymous):

Ok, hold on.

OpenStudy (lexi_loves):

k.

OpenStudy (lexi_loves):

just look at this http://openstudy.com/updates/55b85044e4b0f0c1f1401287

OpenStudy (lexi_loves):

This may help out.... a lot! lol

OpenStudy (anonymous):

When is put the equation into Mathway is said that they are the same, so is sine so sine= −1/√2

OpenStudy (lexi_loves):

Idk. Do you think it is?! :)

OpenStudy (anonymous):

I have no idea.

OpenStudy (lexi_loves):

yes.

OpenStudy (anonymous):

Ok, so tan= -1 sin = √−2/2 cos = √2/2

OpenStudy (lexi_loves):

yea.

OpenStudy (anonymous):

Nice, thank you!!!

OpenStudy (lexi_loves):

but i gtg rn! :) good luck i hope i helped!

OpenStudy (lexi_loves):

your welcome!